OpenStudy (anonymous):

What is the solution of the system? use elimination, i need the steps... x+3y=1 and -5x+4y=-24?

OpenStudy (whpalmer4):

\[x+3y=1\]\[-5x+4y=-24\] Do you see any numbers that could be multiplied with the first equation to give you an equation that has an equal but opposite in sign coefficient with the second equation?

OpenStudy (whpalmer4):

What if you multiply the first one by 5?

OpenStudy (anonymous):

like, so the equations are equal?

OpenStudy (whpalmer4):

Multiply the first equation by 5: \[5*x + 5*3y = 5*1\]\[5x+15y=5\]Add that to the second equation: \[-5x+4y=-24\]------------------\[5x-5x+15y+4y=5-24\]\[0x+19y=-19\]\[19y=-19\]\[y=\] Now you can take that value of \(y\) and put it into either of the original equations and find the value of \(x\).